Calculate the area of a trapezium that has parallel sides of 10 cm and 6 cm respectively and the perpendicular distance of 7 cm between the parallel sides. (5) 


1
Expert's answer
2021-06-29T18:04:54-0400

Given trapezium "ABCD" :

"BE" is height

"AD=10\ cm, BC=6\ cm, BE=7\ cm."


"S_{ABCD}=\dfrac{AD+BC}{2}\cdot BE"

"=\dfrac{10+6}{2}\cdot 7=56(cm^2)"

 The area of a trapezium is 56 cm2.